You are pleased to see that you have been given a 4.48% raise this year.​ However, you read on the Wall Street Journal Web site that inflation over the past year has been 1.88 %

How much better off are you in terms of real purchasing​ power? (Note: Be careful not to round any intermediate steps less than six decimal​ places.) Your real purchasing power is. (Round to two decimal​ places.)
Answer #1

Real purchasing power=(1+nominal rate)/(1+inflation)-1=(1+4.48%)/(1+1.88%)-1=2.552%